Find the dimensions of a rectangle whose perimeter is 46m ans whose area is 126 m^2. (Hint: Let the width be w. Use the perimeter to find the length in terms of w.)
Answer by stanbon(48579)

Perimeter=2l+2w=46
Divide by 2 to get l+w=23
Also, Area=lw=126
Use substitution to solve:
l=23-w
Sustitute to get: (23-w)(w)=126
w^2-23w+126=0
(w-9)(w-14)=0
w=9 or w=14
If w=9, then l=23-9=14
If w=14, then l=23-14=9
So, the dimentions are 9 and 14.
Cheers
Stan H.